# Eliminating Warping Shakes for Unsupervised Online Video Stitching

### v1

@@ Abstract
In this paper, we retarget video stitching to an emerging issue, named warping shake, when extending image stitching to video stitching. It unveils the temporal instability of warped content in non-overlapping regions, despite image stitching having endeavored to preserve the natural structures. Therefore, in most cases, even if the input videos to be stitched are stable, the stitched video will inevitably cause undesired warping shakes and affect the visual experience. To eliminate the shakes, we propose StabStitch to simultaneously realize video stitching and video stabilization in a unified unsupervised learning framework. Starting from the camera paths in video stabilization, we first derive the expression of stitching trajectories in video stitching by elaborately integrating spatial and temporal warps. Then a warp smoothing model is presented to optimize them with a comprehensive consideration regarding content alignment, trajectory smoothness, spatial consistency, and online collaboration. To establish an evaluation benchmark and train the learning framework, we build a video stitching dataset with a rich diversity in camera motions and scenes. Compared with existing stitching solutions, StabStitch exhibits significant superiority in scene robustness and inference speed in addition to stitching and stabilization performance, contributing to a robust and real-time online video stitching system. The code and dataset will be available at https://github.com/nie-lang/StabStitch.
@@ 1 Introduction
Image or video stitching techniques are commonly employed to create panoramic or wide field-of-view (FoV) displays from different viewpoints with limited FoV. Due to their practicality, they are widely applied in autonomous driving [24], video surveillance [38], virtual reality [39], etc.
Image stitching has been studied more extensively and profoundly, which inevitably throws the question of whether existing image stitching solutions can be directly extended to video stitching. Pursuing this line of thought, we initially leverage existing image stitching algorithms [16, 46] to process hand-held camera videos. Subsequently, we observe that although the stitched results for individual frames are remarkably natural, there is obvious content jitter in the non-overlapping regions between temporally consecutive frames. It is also important to note that the jitter does not originate from the inherent characteristics of the source video itself, although these videos are captured by hand-held cameras. In fact, due to the advancements and widespread adoption of video stabilization in both hardware and software nowadays, the source videos obtained from hand-held cameras are typically stable unless deliberately subjected to shaking. For clarity, we define such content jitter as warping shake, which describes the temporal instability of non-overlapping regions induced by temporally non-smooth warps, irrespective of the stability of source videos. Fig. 7 illustrates the occurrence process of warping shakes.
Existing video stitching solutions [48, 52, 11, 31, 17] follow a strong assumption that each source video from freely moving hand-held cameras suffers from heavy and independent shakes. Consequently, every source video necessitates stabilization via warping, contradicting the current prevalent reality that video stabilization technology has already been widely integrated into various portable devices (e.g., cellphones, DV cameras, and UAVs). In addition, these approaches, to jointly optimize video stabilization and stitching, often establish a sophisticated non-linear solving system consisting of various energy terms. To find the optimal parameters, an iterative solving strategy is typically employed. Each iteration involves several steps dedicated to optimizing different parameters separately, resulting in a rather slow inference speed. Moreover, the complicated optimization procedures also impose stringent requirements on the input video quality (e.g., sufficient, accurate, and evenly distributed matching points), making the video stitching systems fragile and not robust in practical applications.
@@ 2.1 Image Stitching
Traditional image stitching methods usually detect keypoints [40] or line segments [55] and then minimize the projective errors to estimate a parameterized warp by aligning these geometric features. To eliminate the parallax misalignment [62], the warp model is extended from global homography transformation [2] to other elastic representations, such as mesh [61], TPS [27], superpixel [25], and triangular facet[26]. Meanwhile, to keep the natural structure of non-overlapping regions, a series of shape-preserving constraints is formulated with the alignment objective. For instance, SPHP [3] and ANAP [30] linearized the homography and slowly changed it to the global similarity to reduce projective distortions; DFW [28], SPW [29], and LPC [16] leveraged line-related consistency to preserve geometric structures; GPS [4] and GES-GPS [7] added a global similarity before stitching multiple images together so that the warp of each image resembles a similar transformation as a whole; etc. Besides, Zhang et al. [63] re-formulated image stitching with regular boundaries by simultaneously optimizing alignment and rectangling [13, 45].
Recently, learning-based image stitching solutions emerged. They feed the entire images into the neural network, encouraging the network to directly predict the corresponding parameterized warp model (e.g., homography [42, 47, 18], multi-homography [51], TPS [46, 21], and optical flow [23, 15]). Compared with traditional methods based on sparse geometric features, these learning-based solutions train the network parameters to adaptively capture semantic features by establishing dense pixel-wise optimization objectives. They show better robustness in various cases, especially in the challenging cases where traditional geometric features are few to detect.
@@ 2.2 Video Stabilization
Traditional video stabilization can be categorized into 3D [32, 35], 2.5D [33, 8], and 2D [41, 10] methods, according to different motion models. The 3D solutions model the camera motions in 3D space or require extra scene structure for stabilization. The structure is either calculated by structure-from-motion (SfM) [32] or acquired from additional hardware, such as a depth camera [35], a gyroscope sensor [20], or a lightfield camera [50]. Given the intensive computational demands of these 3D solutions, 2.5D approaches relax the full 3D requirement to partial 3D information. To this end, some additional 3D constraints are established, such as subspace projection [33] and epipolar geometry [8]. Compared with them, the 2D methods are more efficient with a series of 2D linear transformations (e.g., affine, homography) as camera motions. To deal with large-parallax scenes, spatially varying motion representations are proposed, such as homography mixture [9], mesh [36], vertex profile [34], optical flow [37], etc. Moreover, some special approaches focus on specific input (e.g., selfie [59, 60], 360 [22, 54], and hyperlapse [19] videos).
In contrast, learning-based video stabilization methods directly regress unstable-to-stable transformation from data. Most of them are trained with stable and unstable video pairs acquired by special hardware in a supervised manner [56, 57, 64]. To relieve data dependence, DIFRINT [5] proposed the first unsupervised solution via neighboring frame interpolation. To get a stable interpolated frame, only stable videos are used to train the network. Different from it, DUT [58] established unsupervised constraints for motion estimation and trajectory smoothing, learning video stabilization by watching unstable videos.
@@ 2.3 Video Stitching
Video stitching has received much less attention than image stitching. Early works [17, 49] stitched multiple videos frame-by-frame, and focused on the temporal consistency of stitched frames. But the input videos were captured by cameras fixed on rigs. For hand-held cameras with free and independent motions, there is a significant increase in temporal shakes. To deal with it, videos were first stitched and then stabilized in [12], while [31] did it in an opposite way (e.g., videos were firstly stabilized, and then stitched). Both of them accomplished stitching or stabilization in a separate step. Later, a joint optimization strategy was commonly adopted in [52, 11, 48], where [48] further considered the dynamic foreground by background identification. However, solving such a joint optimization problem regarding stitching and stabilization is fragile and computationally expensive. To this end, we rethink the video stitching problem from the perspective of warping shake and propose the first (to our knowledge) unsupervised online solution for hand-held cameras.
@@ 3.1.1 Temporal Warp:
To obtain camera paths, a temporal warp model is first proposed to represent the temporal motion between consecutive video frames. Different from most video stabilization works [58, 36, 34] that use point correspondences to estimate the warp, we leverage a convolutional neural network to capture the high-level information in inter-frame motions. This alternative proves to be robust across various scenarios, particularly in low-light and low-texture environments. The network structure is similar to the warp network of UDIS++ [46]. As shown in Fig. 2(left), it takes two consecutive target frames as input and outputs the motions of mesh-like control points [1]. Due to the temporal continuity of adjacent frames in the video, the estimated motions are often not significant. Consequently, we replace all global correlation layers [43] in UDIS++ with local correlation layers (i.e., cost volume [53]). To improve the efficiency, we substitute the ResNet50 backbone [14] in UDIS++ with ResNet18 and reduce network parameters accordingly. Following UDIS++, our optimization objective also consists of an alignment term and a distortion term, as described in the following equation:
@@ 3.1.2 Meshflow:
where is the relative transformation from the -th frame to the -th frame. For mesh-like local transformations, it can be further extended as a bundle of spatially varying camera paths for each grid cell :
Considering that our temporal warp model directly predicts the 2D motions of each control point, we chain the motions of each control point temporally as the control point trajectory of the camera for a more straightforward representation:
where is set to zero. Note each control point in is anchored at every vertex in a rigid mesh. Therefore, the previously mentioned camera trajectories, derived from the motions of control points, are also referred to as Meshflow.
@@ 3.2.1 Spatial Warp:
To obtain the stitching trajectories, in addition to the temporal warp model, we also establish a spatial warp model to represent the spatial motion between different video views, as shown in Fig. 2(left). The spatial warp model has a similar network structure to the temporal warp model except that the first local correlation layer is replaced by a global correlation layer [43] to capture long-range matching (usually longer than half of the image width/height). Given the performance of StabStitch heavily relies on the spatial warp, we expect this warp to be as robust as possible, although this network model has been proven to be more robust than traditional methods. To this end, we further introduce a motion consistency term in addition to the basic optimization components of the temporal warp:
@@ 3.2.2 Stitch-Meshflow:
where we define is equal to .
@@ 4 Online Stitching
Existing video stitching methods [48, 52, 11, 31, 17] are offline solutions, which take the whole videos as input, calculate and optimize the complete trajectories from the first frame to the last frame, synthesizing the stitched video after the whole videos are captured and optimized. They typically take a long time (much longer than the duration of the video) to get the final result. Different from them, StabStitch is an online video stitching solution with only one frame latency, which overcomes the above issues (i.e., time-consuming and high-latency). It is more challenging because the trajectories after the current frame are no longer available and real-time inference efficiency is required.
@@ 4.1 Online Smoothing
To achieve this goal, we define a fixed-length sliding window ( frames) to cover previous and current frames, as shown in Fig. 3. Then the local stitching trajectory inside this window is extracted and smoothed according to Sec. 3. Next, the current target frame is re-synthesized using the optimized smooth spatial mesh (Eq. 13). After that, we blend it with the current reference frame to get a stable stitched frame and display the result when the next frame arrives. With this mode and efficient architectures, StabStitch achieves minimal latency with only one frame.
@@
4.2 Dynamic Inference
In addition, we propose a simple dynamic inference strategy that allows for certain adjustments of the balance of distortion and stability. Concretely, for every frame, we first calculate the distortion loss () for the optimized spatial mesh using . Then we strive to reach a balance between the original stitching paths and the smoothed stitching paths by dynamically adjusting the smoothing increment as follows:
where the former component preserves the smoothed paths of overlapping regions and the latter balances the distortion and stability of non-overlapping regions. is the activation function and is to obtain the maximum value. is a threshold parameter ranging from 0 to . For frames with distortion scores exceeding this threshold, an adaptive penalty will be imposed to reduce the degree of distortion at the cost of decreasing stability. One can adjust this threshold to control the degrees of distortion and stability.
@@ 5 Dataset Preparation
We propose a dataset, named StabStitch-D, for the comprehensive video stitching evaluation considering the lack of dedicated datasets for this task. Our dataset comprises over 100 video pairs, consisting of over 100,000 images, with each video lasting from approximately 5 seconds to 100 seconds. To holistically reveal the performance of video stitching methods in various scenarios, we categorize videos into four classes based on their content, including regular (RE), low-texture (LT), low-light (LL), and fast-moving (FM) scenes. In the testing split, 20 video pairs are divided for testing, with 5 videos in each category. Fig. 4 illustrates some examples for each category, where FM is the most challenging case with fast irregular camera movements (rotation or translation). The resolution of each video is resized into for efficient training, and in the testing phase, arbitrary resolutions are supported.
@@ 6.1 Details and Metrics
Please note that in the comparative experiments, we only adopt the alignment score because different methods define different warp representations and camera paths. The last two metrics are used in our ablation study.
@@ 6.2.2 Compared with Video Stitching:
A comprehensive analysis of our inference speed is shown in Tab. 3 with one RTX 4090Ti GPU, where ‘Blending’ represents the average blending. In the example shown in Fig. 5(top), StabStitch only takes about 28.2ms to stitch one frame, yielding a real-time online video stitching system. When stitching a video pair with higher resolution, only the time for warping and blending steps will slightly increase. In contrast, Nie et al’s solution [48] takes over 40 minutes to get such a 26-second stitched video with an Intel i7-9750H 2.60GHz CPU, making it impractical to be applied to online stitching.
@@ 6.3 Ablation Study
The main ablation study is shown in Tab. 4, where ‘Basic Stitching’ (model 1) refers to the spatial warp model without the motion consistency term . With (model 2), the stability is improved. With the warp smoothing model (model 3), both the distortion and stability are significantly optimized at the cost of slight alignment performance, achieving an optimal balance of alignment, distortion, and stability. More experiments can be found in the supplementary.
@@ Appendix 0.A Appendix
Besides, we also provide a demo video (refer to our GitHub page) to demonstrate the stitched results from our StabStitch and other solutions.

### v2

@@ Abstract
In this paper, we retarget video stitching to an emerging issue, named warping shake, when extending image stitching to video stitching. It unveils the temporal instability of warped content in non-overlapping regions, despite image stitching having endeavored to preserve the natural structures. Therefore, in most cases, even if the input videos to be stitched are stable, the stitched video will inevitably cause undesired warping shakes and affect the visual experience. To eliminate the shakes, we propose StabStitch to simultaneously realize video stitching and video stabilization in a unified unsupervised learning framework. Starting from the camera paths in video stabilization, we first derive the expression of stitching trajectories in video stitching by elaborately integrating spatial and temporal warps. Then a warp smoothing model is presented to optimize them with a comprehensive consideration regarding content alignment, trajectory smoothness, spatial consistency, and online collaboration. To establish an evaluation benchmark and train the learning framework, we build a video stitching dataset with a rich diversity in camera motions and scenes. Compared with existing stitching solutions, StabStitch exhibits significant superiority in scene robustness and inference speed in addition to stitching and stabilization performance, contributing to a robust and real-time online video stitching system. The codes and dataset are available at https://github.com/nie-lang/StabStitch.
@@ 1 Introduction
Video stitching techniques are commonly employed to create panoramic or wide field-of-view (FoV) displays from different viewpoints with limited FoV. Due to their practicality, they are widely applied in autonomous driving [24], video surveillance [38], virtual reality [39], etc. Our work lies in the most common and challenging case of video stitching with hand-held cameras. It does not require camera poses, motion trajectories, or temporal synchronization. It merges multiple videos, whether from multiple cameras or a single camera capturing multiple videos, to create a more immersive representation of the captured scene. Moreover, it transforms video production into an enjoyable and collaborative endeavor among a group of individuals.
Compared with video stitching, image stitching has been studied more extensively and profoundly, which inevitably throws the question of whether existing image stitching solutions can be directly extended to video stitching. Pursuing this line of thought, we initially leverage existing image stitching algorithms [16][47] to process hand-held camera videos. Subsequently, we observe that although the stitched results for individual frames are remarkably natural, there is obvious content jitter in the non-overlapping regions between temporally consecutive frames. It is also important to note that the jitter does not originate from the inherent characteristics of the source video itself, although these videos are captured by hand-held cameras. In fact, due to the advancements and widespread adoption of video stabilization in both hardware and software nowadays, the source videos obtained from hand-held cameras are typically stable unless deliberately subjected to shaking. For clarity, we define such content jitter as warping shake, which describes the temporal instability of non-overlapping regions induced by temporally non-smooth warps, irrespective of the stability of source videos. Fig. 1 illustrates the occurrence process of warping shakes.
Existing video stitching solutions [49][53][11][31][17] follow a strong assumption that each source video from freely moving hand-held cameras suffers from heavy and independent shakes. Consequently, every source video necessitates stabilization via warping, contradicting the current prevalent reality that video stabilization technology has already been widely integrated into various portable devices (e.g., cellphones, DV cameras, and UAVs). In addition, these approaches, to jointly optimize video stabilization and stitching, often establish a sophisticated non-linear solving system consisting of various energy terms. To find the optimal parameters, an iterative solving strategy is typically employed. Each iteration involves several steps dedicated to optimizing different parameters separately, resulting in a rather slow inference speed. Moreover, the complicated optimization procedures also impose stringent requirements on the input video quality (e.g., sufficient, accurate, and evenly distributed matching points), making the video stitching systems fragile and not robust in practical applications.
@@ 2.1 Image Stitching
Traditional image stitching methods usually detect keypoints [40] or line segments [56] and then minimize the projective errors to estimate a parameterized warp by aligning these geometric features. To eliminate the parallax misalignment [63], the warp model is extended from global homography transformation [2] to other elastic representations, such as mesh [62], TPS [27], superpixel [25], and triangular facet[26]. Meanwhile, to keep the natural structure of non-overlapping regions, a series of shape-preserving constraints is formulated with the alignment objective. For instance, SPHP [3] and ANAP [30] linearized the homography and slowly changed it to the global similarity to reduce projective distortions; DFW [28], SPW [29], and LPC [16] leveraged line-related consistency to preserve geometric structures; GSP [4] and GES-GSP [7] added a global similarity before stitching multiple images together so that the warp of each image resembles a similar transformation as a whole; etc. Besides, Zhang et al. [64] re-formulated image stitching with regular boundaries by simultaneously optimizing alignment and rectangling [13][46].
Recently, learning-based image stitching solutions emerged. They feed the entire images into the neural network, encouraging the network to directly predict the corresponding parameterized warp model (e.g., homography [43][48][18], multi-homography [52], TPS [47][21][65], and optical flow [23][15]). Compared with traditional methods based on sparse geometric features, these learning-based solutions train the network parameters to adaptively capture semantic features by establishing dense pixel-wise optimization objectives. They show better robustness in various cases, especially in the challenging cases where traditional geometric features are few to detect.
@@ 2.2 Video Stabilization
Traditional video stabilization can be categorized into 3D [32][35], 2.5D [33][8], and 2D [42][10][41] methods, according to different motion models. The 3D solutions model the camera motions in 3D space or require extra scene structure for stabilization. The structure is either calculated by structure-from-motion (SfM) [32] or acquired from additional hardware, such as a depth camera [35], a gyroscope sensor [20], or a lightfield camera [51]. Given the intensive computational demands of these 3D solutions, 2.5D approaches relax the full 3D requirement to partial 3D information. To this end, some additional 3D constraints are established, such as subspace projection [33] and epipolar geometry [8]. Compared with them, the 2D methods are more efficient with a series of 2D linear transformations (e.g., affine, homography) as camera motions. To deal with large-parallax scenes, spatially varying motion representations are proposed, such as homography mixture [9], mesh [36], vertex profile [34], optical flow [37], etc. Moreover, some special approaches focus on specific input (e.g., selfie [60][61], 360 [22][55], and hyperlapse [19] videos).
In contrast, learning-based video stabilization methods directly regress unstable-to-stable transformation from data. Most of them are trained with stable and unstable video pairs acquired by special hardware in a supervised manner [57][58][66]. To relieve data dependence, DIFRINT [5] proposed the first unsupervised solution via neighboring frame interpolation. To get a stable interpolated frame, only stable videos are used to train the network. Different from it, DUT [59] established unsupervised constraints for motion estimation and trajectory smoothing, learning video stabilization by watching unstable videos.
@@ 2.3 Video Stitching
Video stitching has received much less attention than image stitching. Early works [17][50] stitched multiple videos frame-by-frame, and focused on the temporal consistency of stitched frames. But the input videos were captured by cameras fixed on rigs. For hand-held cameras with free and independent motions, there is a significant increase in temporal shakes. To deal with it, videos were first stitched and then stabilized in [12], while [31] did it in an opposite way (e.g., videos were firstly stabilized, and then stitched). Both of them accomplished stitching or stabilization in a separate step. Later, a joint optimization strategy was commonly adopted in [53][11][49], where [49] further considered the dynamic foreground by background identification. However, solving such a joint optimization problem regarding stitching and stabilization is fragile and computationally expensive. To this end, we rethink the video stitching problem from the perspective of warping shake and propose the first (to our knowledge) unsupervised online solution for hand-held cameras.
@@ 3.1.1 Temporal Warp:
To obtain camera paths, a temporal warp model is first proposed to represent the temporal motion between consecutive video frames. Different from most video stabilization works [59][36][34] that use point correspondences to estimate the warp, we leverage a convolutional neural network to capture the high-level information in inter-frame motions. This alternative proves to be robust across various scenarios, particularly in low-light and low-texture environments. The network structure is similar to the warp network of UDIS++ [47]. As shown in Fig. 2(left), it takes two consecutive target frames as input and outputs the motions of mesh-like control points [1]. Due to the temporal continuity of adjacent frames in the video, the estimated motions are often not significant. Consequently, we replace all global correlation layers [44] in UDIS++ with local correlation layers (i.e., cost volume [54]). To improve the efficiency, we substitute the ResNet50 backbone [14] in UDIS++ with ResNet18 and reduce network parameters accordingly. Following UDIS++, our optimization objective also consists of an alignment term and a distortion term, as described in the following equation:
@@ 3.1.2 Meshflow:
where is the relative transformation from the -th frame to the -th frame. Considering that our temporal warp model directly predicts the 2D motions of each control point, we adopt the motion representation of vertex files like MeshFlow [34]. Particularly, we chain the motions of each control point temporally as the control point trajectory for a more straightforward representation:
where is set to zero. Note each control point in is anchored at every vertex in a rigid mesh.
@@ 3.2.1 Spatial Warp:
To obtain the stitching trajectories, in addition to the temporal warp model, we also establish a spatial warp model to represent the spatial motion between different video views, as shown in Fig. 2(left). The spatial warp model has a similar network structure to the temporal warp model except that the first local correlation layer is replaced by a global correlation layer [44] to capture long-range matching (usually longer than half of the image width/height). Considering the significance of spatial warping stability in video stitching, we expect this warp to be as robust as possible, although this network model has been proven to be more robust than traditional methods. To this end, we further introduce a motion consistency term in addition to the basic optimization components of the temporal warp:
@@ 3.2.2 Stitch-Meshflow:
where we define is an all-zero array.
@@ 4 Online Stitching
Existing video stitching methods [49][53][11][31][17] are offline solutions, which smooth the trajectories after the videos are completely captured. Different from them, StabStitch is an online video stitching solution. In our case, the frames after the current frame are no longer available and real-time inference is required.
@@ 4.1 Online Smoothing
To achieve this goal, we define a fixed-length sliding window ( frames) to cover previous and current frames, as shown in Fig. 3. Then the local stitching trajectory inside this window is extracted and smoothed according to Sec. 3. Next, the current target frame is re-synthesized using the optimized smooth spatial mesh (Eq. 12). Finally, we blend it with the current reference frame to get a stable stitched frame and display the result when the next frame arrives. With this mode and efficient architectures, StabStitch achieves minimal latency with only one frame.
@@
4.2 Offline and Online Inference
Offline smoothing takes the whole trajectories as input, outputs the optimized whole trajectories, and then renders all the video frames. It carries on smoothing after receiving whole input videos and can be regarded as a special online case in which the sliding window covers whole videos. By contrast, online smoothing takes local trajectories as input, outputs the optimized local trajectories, and then renders the last frame in the sliding window. The online process smoothes current paths without using any future frames.
@@ 5 Dataset Preparation
We establish a dataset, named StabStitch-D, for the comprehensive video stitching evaluation considering the lack of dedicated datasets for this task. Our dataset comprises over 100 video pairs, consisting of over 100,000 images, with each video lasting from approximately 5 seconds to 35 seconds. To holistically reveal the performance of video stitching methods in various scenarios, we categorize videos into four classes based on their content, including regular (RE), low-texture (LT), low-light (LL), and fast-moving (FM) scenes. In the testing split, 20 video pairs are divided for testing, with 5 videos in each category. Fig. 4 illustrates some examples for each category, where FM is the most challenging case with fast irregular camera movements (rotation or translation). The resolution of each video is resized into for efficient training, and in the testing phase, arbitrary resolutions are supported.
@@ 6.1 Details and Metrics
To quantitatively evaluate the proposed method, we suggest three metrics including alignment score, distortion score, and stability score. Limited by space, we moved the related metric details to the supplementary.
@@ 6.2.2 Compared with Video Stitching:
A comprehensive analysis of our inference speed is shown in Tab. 3 with one RTX 4090Ti GPU, where ‘Blending’ represents the average blending. In the example shown in Fig. 5(top), StabStitch only takes about 28.2ms to stitch one frame, yielding a real-time online video stitching system. When stitching a video pair with higher resolution, only the time for warping and blending steps will slightly increase. In contrast, Nie et al.’s solution [49] takes over 40 minutes to get such a 26-second stitched video with an Intel i7-9750H 2.60GHz CPU, making it impractical to be applied to online stitching.
@@ 6.3 Ablation Study
The main ablation study is shown in Tab. 4, where ‘Basic Stitching’ (model 1) refers to the spatial warp model without the motion consistency term . With (model 2), the stability is improved. With the warp smoothing model (model 3), both the distortion and stability are significantly optimized at the cost of slight alignment performance, achieving an optimal balance of alignment, distortion, and stability. More experiments can be found in the supplementary materials.
@@ 7 Conclusions
Acknowledgments: This work was supported by the National Natural Science Foundation of China (No. 62172032), and Zhejiang Province Basic Public Welfare Research Program (No. LGG22F020009).
@@ Appendix 0.A Appendix
For clarity, we summarize a part of notations and their corresponding meanings in Table 5.
Besides, we also provide a supplementary video. Please refer to https://www.youtube.com/watch?v=03kGEZJHxzI for the stitched videos from different methods.
@@ 0.B.1 Structure Difference
| Notation | Meaning | Example |
| The relative time in a sliding window. | e.g., C(t), m(t), S(t) | |
| Subscript | The control point index. | e.g., , , |
| Superscript , | From the temporal or spatial model. | e.g., , |
| Superscript | The absolute time ranging from to the last frame. | e.g., , |
| Hat | The optimized mesh or trajectory. | e.g., , |
| The rigid and regular initial mesh (predefined). | ||
| The -th reference/target frame. | ||
| The spatial/temporal/smooth warping models. | ||
| The TPS transformation from to . |